Jiminy's Cricket Farm issued a bond with 30 years to maturity and a semiannual coupon rate of 6 percent 4 years ago. The bond currently sells for 105 percent of its face value. The company's tax rate is 23 percent. The book value of the debt issue is $60 million. In addition, the company has a second debt issue on the market, a zero coupon bond with 8 years left to maturity, the book value of this issue is $35 million, and the bonds sell for 67 percent of par. 

What is the company's total book value of debt? 

What is the company's total market value of debt?

What is your best estimate of the aftertax cost of debt?
Answer #1
  1. Book Value of Debt = 60 million + 35 Million = 95 million

    b. Market Value of Debt = 105%*60+67%*35 = 86.45 million

    c. YTM of first bond is calculated as follows

    Let's assume par value of single bond = 1000
    Coupon= 6%*1000/2 =30
    Price = 105%*1000 = 1050
    Number of Periods = Number of Years Remaining*2 =(30-4)*2 = 52
    YTM using excel formula =2*RATE(52,30,-1050,1000) = 5.6315%


    Price of second bond = 67%*1000 = 670
    Number of periods = 8
    YTM of second bond = (1000/670)^(1/8)-1 = 5.1334%

    Cost of debt = Cost of Debt * Weight of market value of debt + Cost of zero coupon debt* Weight of market value of zero coupon = 5.6315%*(105%*60/86.45)+5.1334%*(67%*35/86.45) = 5.4964%

    After tax cost of Debt = 5.4964%*(1-Tax rate) = 5.4964%*(1-23%) = 4.23%

    ​​​​​​​